OpenStudy (anonymous):

Solve: 10^x=1/3 Oh pleaseeeeeeeeeeeeeeeeeee help!!!!!!!!

OpenStudy (anonymous):

log10(1/3) = -Log10(3) =

OpenStudy (anonymous):

By Log10, I mean the base 10 log, on your calculator as LOG

OpenStudy (vijay):

\[-\log (3) / [ \log 2 + \log 5]\]

OpenStudy (anonymous):

-.4771212

OpenStudy (anonymous):

thanks everyone that helped :) ♥

OpenStudy (anonymous):

Thank YOU for the help. :)

OpenStudy (anonymous):

xDDD welcome :)